Coca-Cola, Inc. (KO) expects to pay a dividend of $2.10 per share next year. Historically their...

  1. Coca-Cola, Inc. (KO) expects to pay a dividend of $2.10 per share next year. Historically their dividends have been growing at a constant rate of about 5% per year. Based on the risk of KO, investor’s require a 7.5% return. Estimate the current stock price for Coca-Cola.

    A.

    $76.35

    B.

    $84.00

    C.

    $91.45

    D.

    $103.21

    E.

    cannot be determined

2.

  1. Coca Cola is evaluating an investment in a new syrup production machine. The initial investment in the project is $110,000. It has been estimated that annual cash inflows of $40,000 will be generated by the machine for the next 5 years. The opportunity cost of the project is estimated to be 6%. Calculate the Net Present Value (NPV) of the project.

    A.

    $58,495

    B.

    $90,000

    C.

    $43,218

    D.

    $34,875

    E.

    $64,755

3.

  1. Coca Cola is evaluating an investment in a new syrup production machine. The initial investment in the project is $110,000. It has been estimated that annual cash inflows of $40,000 will be generated by the machine for the next 5 years. The opportunity cost of the project is estimated to be 6%. Calculate the Internal Rate of Return (IRR) of the

    project.

    A.

    6%

    B.

    12.7%

    C.

    23.9%

    D.

    19.6%

    E.

    32.8%

4.

  1. Coca Cola is evaluating an investment in a new syrup production machine. The initial investment in the project is $110,000. It has been estimated that annual cash inflows of $40,000 will be generated by the machine for the next 5 years. The opportunity cost of the project is estimated to be 6%. Calculate the Payback Period (PB) of the project.

    A.

    5 years

    B.

    2.5 years

    C.

    2.75 years

    D.

    3.75 years

    E.

    3.5 years
Answer #1
1) B. $         84.00
Working:
Current Stock Price = D1/(Ke-g) Where,
= 2.10/(0.0750-0.05) D1 = $       2.10
= $               84.00 Ke = 7.50%
g = 5%
2) A. $       58,495
Working;
Year Cash inflow Discount factor Present Value
a b c=1.06^-a d=b*c
0 $       -1,10,000           1.0000 $       -1,10,000
1 $             40,000           0.9434 $             37,736
2 $             40,000           0.8900 $             35,600
3 $             40,000           0.8396 $             33,585
4 $             40,000           0.7921 $             31,684
5 $             40,000           0.7473 $             29,890
NPV $             58,495
3) C. 23.9%
Working:
Year Cash flow
0 $ -1,10,000
1 $       40,000
2 $       40,000
3 $       40,000
4 $       40,000
5 $       40,000
IRR =irr(C30:C35)
23.9%
C30:C35 is the range of cash flow from year 0 to 5.
4) C. 2.75 Years
Working:
Year Cash flow Cumulative cash flow
0 $ -1,10,000 $       -1,10,000
1 $       40,000 $           -70,000
2 $       40,000 $           -30,000
3 $       40,000 $             10,000
4 $       40,000 $             50,000
5 $       40,000 $             90,000
Payback period = 2+(30000/40000)
= 2.75 Years